### 15 Jan 2024 v2

# Addressing Unmeasured Confounders in Cox Proportional Hazards Models Using Nonparametric Bayesian Approaches

Abstract
In observational studies, unmeasured confounders present a crucial challenge in accurately estimating desired causal effects. To calculate the hazard ratio (HR) in Cox proportional hazard models, which are relevant for time-to-event outcomes, methods such as Two-Stage Residual Inclusion and Limited Information Maximum Likelihood are typically employed. However, these methods raise concerns, including the potential for biased HR estimates and issues with parameter identification. This manuscript introduces a novel nonparametric Bayesian method designed to estimate an unbiased HR, addressing concerns related to parameter identification. Our proposed method consists of two phases: 1) detecting clusters based on the likelihood of the exposure variable, and 2) estimating the hazard ratio within each cluster. Although it is implicitly assumed that unmeasured confounders affect outcomes through cluster effects, our algorithm is well-suited for such data structures.
Keywords: frailty model, general Bayes, Mendelian randomization, nonparametric Bayes, unmeasured confounders
1 Introduction
In observational studies, unmeasured confounders denote situations where certain confounders are not observed. In such cases, instrumental variable (IV) methods are commonly employed. In the fields of biometrics and related areas, several theoretical results and applications of IV methods have emerged in recent years. Mendelian Randomization (MR), an IV method that typically uses single nucleotide polymorphisms (SNPs) as IVs, is one such example.
Time-to-event outcomes are frequently encountered in biometrics and related fields. The Cox Proportional Hazards Model (CPHM; Cox, 1972) is commonly used for analyzing these outcomes while adjusting for confounders. However, obtaining a consistent estimator can be challenging when some confounders are unobserved (Gail et al., 1984). Martínez-Camblor et al. (2019) employed the Two-Stage Residual Inclusion (2SRI) approach, an instrumental variable technique (Terza et al., 2008), using the frailty model (Nielsen et al., 1992) to address unobserved variability. They explored useful properties of their estimator under continuous exposure conditions. Nonetheless, 2SRI may produce biased causal effects in certain scenarios (Basu et al., 2017; Wan et al., 2018; Orihara et al., 2023). Basu et al. (2017) and Orihara et al. (2023) demonstrated that a full-likelihood approach yields more accurate estimates than 2SRI (also discussed in Section 5 of Burgess et al., 2017). The limited information maximum likelihood (LIML) estimator employs a full-likelihood approach and shares similarity with 2SRI. Orihara (2022) proposed a LIML-based hazard ratio estimator that is applicable not only in continuous exposure scenarios but also in binary exposure contexts. Unfortunately, this method requires the specification of a variance parameter for unmeasured confounders prior to the analysis. While they mentioned that this parameter could be treated as a sensitivity parameter, that can derive similar interpretation as the E-value (VanderWeele and Ding, 2018), it’s more common in analyses to estimate point estimates and confidence bands for interpretation of results.
In this manuscript, we introduce a novel Bayesian-based hazard ratio estimator. The model construction of this estimator is conceptually similar to the LIML-based estimator. Specifically, it employs nonparametric Bayesian techniques with cluster construction to account for unmeasured confounders. To estimate the hazard ratio accurately, our approach involves two phases: 1) detecting clusters based on the likelihood for the exposure variable, and 2) estimating the hazard ratio within each cluster. The first step parallels the spirit of propensity score subclassification (Rosenbaum and Rubin, 1984; Imai and van Dyk, 2004), as each cluster is considered to have a similar value of the (generalized) propensity score. To accurately construct the second step within the Bayesian paradigm, we utilize general Bayes methods (Bissiri et al., 2016) without the need for baseline hazard modeling. While there are some related topics using Bayesian methods, our unique contribution lies in the application of nonparametric Bayesian methods to the CPHM.
The remainder of this paper is organized as follows: Section 2 introduces the model and assumptions used in this study. Section 3 details the proposed Bayesian methods and their algorithms, including an in-depth consideration of the proposed procedures. Section 4 validates the properties of the proposed method through simulations.
2 Preliminaries
[…]
$A_{k}$ $=\alpha_{k}+\boldsymbol{z}^{\top}\boldsymbol{\alpha}_{z}+\boldsymbol{x}^{\top}\boldsymbol{\alpha}_{x}+\tilde{\varepsilon}_{k}=\boldsymbol{z}^{\top}\boldsymbol{\alpha}_{z}+\boldsymbol{x}^{\top}\boldsymbol{\alpha}_{x}+\varepsilon_{k},$ (2.1)
$\lambda_{k}(t)$ $=\lambda_{0k}(t)\exp\left\{\tilde{\boldsymbol{x}}_{t}^{\top}\boldsymbol{\beta}\right\}$ (2.2)
where $\tilde{\varepsilon}_{k}$ and $\varepsilon_{k}:=\alpha_{k}+\tilde{\varepsilon}_{k}$ are error terms with ${\rm E}[\tilde{\varepsilon}_{k}]=0$ and $Var(\tilde{\varepsilon}_{k})=\sigma^{2}_{k}$, and $\lambda_{0k}(t)$ indicates the baseline hazard and $\tilde{\boldsymbol{x}}_{t}$ represents the vector of explanatory variables for the outcome model. For instance, $\tilde{\boldsymbol{x}}_{t}$ includes $w$ and $\boldsymbol{x}$; $\tilde{\boldsymbol{x}}_{t}^{\top}=\left(a,\boldsymbol{x}^{\top}\right)$ and $\boldsymbol{\beta}^{\top}=\left(\beta_{a},\boldsymbol{\beta}_{x}^{\top}\right)$. In this context, we are interested in the treatment effect $\beta_{a}$. The variable $k\in{1,2,\dots,K}$ represents an unknown cluster to which a subject belongs, and we do not know the total number of clusters, denoted as $K$. This uncertainty presents a challenge in estimating the parameters. Note that the interaction term between the exposure and confounders can be included in the model (2.2).
In the models (2.1) and (2.2), each $k$ can be interpreted as a cluster effect resulting from unmeasured confounders. For instance, consider a scenario where a larger $k$ corresponds to a larger value of $\alpha_{k}$ and $\lambda_{0k}(t)$ in the early time period. In such cases, subjects with larger exposure values tend to experience the event of interest more rapidly. This suggests that the treatment effect $\beta_{w}$ could be biased if the cluster effects are not considered. This bias is a well-known phenomenon referred to as ‘confounding bias.’ Therefore, to accurately estimate the treatment effect $\beta_{w}$, it is essential to account for the cluster effect, i.e., the impact of unmeasured confounders (see Figure 1).
Figure 1: Image of the impact of cluster effects as confounding bias
Confounders that may have different distributions across each cluster need not necessarily be included in model (2.1) (i.e., $\boldsymbol{\alpha}_{x}\equiv\boldsymbol{0}$). However, IVs that are not affected by cluster effects are crucial for accurately estimating $\alpha_{k}$s; they play an important role in identifying each cluster. For example, if the distribution of a single IV is proportional to the cluster number $k$ (i.e., a larger $k$ corresponds to a larger value of $Z$), then $\alpha_{k}$s will not be effective in distinguishing between clusters; in such cases, $\alpha_{k}\equiv\alpha$ for all $k\in\{1,2,\dots,K\}$. The conditions for satisfying the IVs correspond to parts of the IV assumptions.
The model (2.2) can be considered as the frailty model (Nielsen et al., 1992) since the baseline hazards functions are different in each cluster. In other words, the outcome value is affected by only through cluster effects by the baseline hazard function; not through the treatment effects. This is corresponded to the exclusion restriction assumption in the IV assumptions. Therefore, the models (2.1) and (2.2) satisfy the three instrumental variable assumptions.
2.1 Likelihoods to Construct the Bayesian-based Estimation
In addition to the assumptions for (2.1), we assume that $\varepsilon_{k}$ follows a normal distribution: $\varepsilon_{k}\sim N(\alpha_{k},\sigma_{k}^{2})$. This assumption indicates that the error term $\varepsilon_{k}$ does not necessarily need to be consistent within each cluster; heterogeneity is permitted. To effectively identify each cluster, it is essential to include as many predictors as possible in model (2.1). As obviously, the likelihood for the exposure variable can be expressed as follows:
$L(\boldsymbol{\theta})=\prod_{k=1}^{K}L(\boldsymbol{\theta}_{k})=\prod_{k=1}^{K}\prod_{i=1}^{n_{k}}\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left\{-\frac{(A_{ik}-\boldsymbol{w}_{ik}^{\top}\boldsymbol{\alpha}_{k})^{2}}{2\sigma_{k}^{2}}\right\},$ (2.3)
where $\boldsymbol{w}_{ik}^{\top}\boldsymbol{\alpha}_{k}=\alpha_{k}+\boldsymbol{z}^{\top}\boldsymbol{\alpha}_{z}+\boldsymbol{x}^{\top}\boldsymbol{\alpha}_{x}$, and $\boldsymbol{\theta}_{k}:=(\boldsymbol{\alpha}_{k}^{\top},\sigma_{k})^{\top}$. The model is similar to, but not identical to, the likelihood assumed in Orihara (2022). In Orihara (2022), it is assumed that all subjects share the same error distribution (i.e., $\boldsymbol{\theta}_{k}\equiv\boldsymbol{\theta}$ for all $k\in\{1,\dots,K\}$), and that $\sigma$ is a fixed value. However, the assumptions for the likelihood (2.3) are somewhat more relaxed. Therefore, the likelihood can be seen as a direct extension of the LIML-based methods.
For model (2.2), we use a partial likelihood for each cluster instead of a full likelihood:
$pl_{k}(\boldsymbol{\beta})=\prod_{i=1}^{n_{k}}\left(\frac{\exp\left\{\tilde{\boldsymbol{x}}_{i,t}^{\top}\boldsymbol{\beta}\right\}}{\sum_{\ell\in R_{k}(t_{ik})}\exp\left\{\tilde{\boldsymbol{x}}_{\ell,t}^{\top}\boldsymbol{\beta}\right\}}\right)^{\delta_{i}},$ (2.4)
where $\delta_{i}$ is the indicator whether an event is observed ($\delta_{i}=1$) or not ($\delta_{i}=0$), and $R_{k}(t)$ denotes the risk set of the cluster $k$ at $t$. As detailed in the following section, constructing a Bayesian-based estimation procedure typically requires a full likelihood (Kalbfleisch and Prentice, 2011) and assuming a prior distribution for the baseline hazard function (Sinha et al., 2003). However, since we are generally not interested in the baseline hazard function, we develop a Bayesian procedure using the partial likelihood (2.4).
3 Nonparametric Bayesian Estimating Procedure
To estimate the target parameter $\boldsymbol{\beta}$, we propose a novel nonparametric Bayesian estimation method. To implement this, we consider not only the likelihoods (2.3) and (2.4) but also prior distributions for the parameters. For $\boldsymbol{\beta}$, the prior distribution is not typically restrictive; we can employ a distribution that reflects prior information before data collection, use noninformative improper prior distributions, and so on.
However, for $\boldsymbol{\theta}_{k}$, we assume the Dirichlet process prior in this manuscript (Escobar and West, 1998; Conley et al., 2008; Müller et al., 2015). Specifically, we assume that $\boldsymbol{\theta}_{k}$ follows hierarchical models: $\boldsymbol{\theta}_{k}|G\sim G(\boldsymbol{\theta}_{k}),\ G(\cdot)\sim\mathcal{D}(G|G_{0},\gamma)$, where $\mathcal{D}$ represents the Dirichlet process prior, and $G$ and $G_{0}$ are referred to as the ‘prior distribution’ and ‘base distribution’, respectively. When there are cluster structures within the data, the Dirichlet process prior can estimate parameters with clustering, without requiring any information about the clusters or their sizes. The size of each cluster is controlled by $\gamma$, known as the ‘precision parameter’.
To construct a Bayesian procedure, it is commonly assumed that data information is reflected in the parameter of interest through a specific likelihood function. However, as mentioned in the previous section, we assume a partial likelihood for the parameter $\boldsymbol{\beta}$. As discussed in Bissiri et al. (2016), a loss function that includes the partial likelihood for the CPHM can be justified from a Bayesian perspective. In other words, the partial likelihood can be applied in the same manner as ordinary likelihoods.
3.1 Proposed Estimating Procedure
Summarizing the above discussions, we propose the following sampling algorithm to sample from the posterior distribution. Note that $\pi(\cdot)$ and $\pi(\cdot|\cdot)$ represent the prior and posterior distributions, respectively. Also, let $O:=(O_{1},\dots,O_{n})$ denote the set of observed data (i.e., $O_{i}=(t_{i},\delta_{i},a_{i},\boldsymbol{x}_{i},\boldsymbol{z}_{i})$).
Posterior for $\boldsymbol{\alpha}_{z}$ and $\boldsymbol{\alpha}_{x}$
Sampling from the following posterior distribution:
$\pi(\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x}|O,\boldsymbol{\beta},\alpha_{k},\sigma_{k})$ $\propto L(\boldsymbol{\theta})\pi(\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x})=\prod_{k=1}^{K}\prod_{i=1}^{n_{k}}\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left\{-\frac{(A_{ik}-\boldsymbol{w}_{ik}^{\top}\boldsymbol{\alpha}_{k})^{2}}{2\sigma_{k}^{2}}\right\}\pi(\boldsymbol{\alpha}_{z})\pi(\boldsymbol{\alpha}_{x})$ (3.1)
When $\pi(\boldsymbol{\alpha}_{z})$ and $\pi(\boldsymbol{\alpha}_{x})$ are normal distributions, (3.1) also becomes a normal distribution. However, we use Stan (Carpenter et al., 2017) for sampling, as we assume a more complex prior distribution in the following simulation experiments.
Posterior for $\boldsymbol{\beta}$
We would like to sample from the following posterior distribution:
$\pi(\boldsymbol{\beta}|O,\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x},\alpha_{k},\sigma_{k})$ $\propto\prod_{k=1}^{K}\prod_{i=1}^{n_{k}}f(t_{ik},\delta_{ik}|a_{ik},\boldsymbol{x}_{ik},\boldsymbol{\beta})\pi(\boldsymbol{\beta})$ (3.2)
As mentioned previously, when sampling from (3.2), it is not necessary to take into account the baseline hazard function $\lambda_{0k}(t)$. Specifically, by utilizing the likelihood (2.4), we employ the following sampling distribution:
$\pi(\boldsymbol{\beta}|O,\boldsymbol{\beta},\alpha_{k},\sigma_{k})$ $\propto\prod_{k=1}^{K}pl_{k}(\boldsymbol{\beta})\pi(\boldsymbol{\beta})=\prod_{k=1}^{K}\prod_{i=1}^{n_{k}}\left(\frac{\exp\left\{\tilde{\boldsymbol{x}}_{i,t}^{\top}\boldsymbol{\beta}\right\}}{\sum_{\ell\in R_{k}(t_{ik})}\exp\left\{\tilde{\boldsymbol{x}}_{\ell,t}^{\top}\boldsymbol{\beta}\right\}}\right)^{\delta_{i}}\pi(\boldsymbol{\beta})$ (3.3)
We also use Stan for sampling from (3.3).
Posterior for $\alpha_{k}$ and $\sigma_{k}$
To achieve sampling from $\pi(\boldsymbol{\theta}|O,\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x},\boldsymbol{\beta})$, we construct a sampling techniques using a Polya urn shceme representation (see Escobar and West, 1998; Conley et al., 2008). Specifically, for each iteration of Gibbs sampling, $\boldsymbol{\theta}_{k}$s are sampled from the following distribution:
$\boldsymbol{\theta}_{i}|\boldsymbol{\theta}_{-i},O,\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x},G_{0},\gamma\sim\left\{\begin{array}[]{l}{\rm with\ probability}\ q_{0}\ {\rm draw\ from}\ \pi(\boldsymbol{\theta}_{i}|O_{i},\lambda)\\ {\rm with\ probability}\ q_{k^{*}}\ {\rm draw\ from}\ \delta_{k^{*}}\end{array}\right.,$ (3.6)
where
$\pi(\boldsymbol{\theta}_{i}|O_{i},\lambda)$ $\propto\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left\{-\frac{(A_{ik}-\boldsymbol{w}_{ik}^{\top}\boldsymbol{\alpha}_{k})^{2}}{2\sigma_{k}^{2}}\right\}G_{0}(\boldsymbol{\theta}_{i}),$
$q_{0}$ $\propto\lambda\int\frac{1}{\sqrt{2\pi\sigma_{k}^{2}}}\exp\left\{-\frac{(A_{ik}-\boldsymbol{w}_{ik}^{\top}\boldsymbol{\alpha}_{k})^{2}}{2\sigma_{k}^{2}}\right\}G_{0}(d\boldsymbol{\theta}_{i}),$
$q_{k^{*}}$ $\propto\frac{1}{\sqrt{2\pi\sigma_{k^{*}}^{2}}}\exp\left\{-\frac{(A_{ik}-\boldsymbol{w}_{ik^{*}}^{\top}\boldsymbol{\alpha}_{k^{*}})^{2}}{2\sigma_{k^{*}}^{2}}\right\}$
with $q_{0}+\sum_{k^{*}}n_{k^{*}}q_{k^{*}}=1$, where $n_{k^{*}}$ is the number of subjects in cluster $k^{*}$, and $\delta_{k^{*}}$ is the Dirac delta function at $\boldsymbol{\theta}_{k^{*}}$. By iterating the sampling from $\boldsymbol{\theta}_{i}|\boldsymbol{\theta}_{-i},O,\boldsymbol{\alpha}_{z},\boldsymbol{\alpha}_{x},G_{0},\gamma$, clusters are automatically created, adjusted by the prior $\lambda$. In the simulation experiments, we assume $G_{0}$ follows a normal distribution to simplify the calculation.
From the above sampling scheme, we can construct marginal posterior distribution for $\boldsymbol{\alpha}_{z}$, $\boldsymbol{\alpha}_{x}$, $\boldsymbol{\beta}$, $\alpha_{k}$, and $\sigma_{k}$. In the following simulation experiments, we confirm performance of our proposed estimating procedure.
4 Simulation Experiments
In this section, we evaluate the performance of our proposed methods compared to several competitors, which will be explained later, using simulation data examples. The number of iterations for all simulation examples was set to 300.
4.1 Data-generating Mechanism
First, we will describe the Data-Generating Mechanism (DGM) used in this simulation. We consider one unmeasured confounder, denoted as $U_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Bin(2,0.5)$. This unmeasured condounder derives the cluster effect which we cannot identify it directly. Also, we consider one instrumental variable, denoted as $Z_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,5^{2})$. Using these variables, the exposure variable is defined by $A|z_{i},u_{i}\sim N(10+z_{i}+4\times(u_{i}-0.5),1)$. Here, $\varepsilon$ adjusts the relationship between the assignment and the confounders, influencing the variance of the propensity score. The outcome model is specified as $\lambda(t_{i}|a_{i},u_{i})=\lambda_{0}(t_{i},u_{i})\exp\left\{-0.05\times a_{i}\right\}$, where $\lambda_{0}(t_{i},u_{i})=0.1\times\left(u_{i}+1\right)$. Finally, the uniform censoring is included in 10%.
4.2 Simulation Results
The crude mean and standard deviation for the exposure variable are 12.28 (3.99). The histogram for the exposure variable and the Kaplan-Meier curve for the outcome variable are shown in Figures 3 and 4, respectively. No obvious cluster effects are observed. However, it is necessary to account for hidden, unmeasured confounder effects to estimate an unbiased hazard ratio.
The descriptive statistics for the estimated hazard ratios are summarized in Table 1 and Figure 2. Our proposed estimator demonstrates a degree of accuracy in comparison with the Cox regression estimator, which is efficient when the unmeasured confounder is used. In contrast, the naive Cox regression estimator shows a significant bias. Consequently, it is anticipated that the proposed method will perform effectively, even in the presence of some unmeasured confounders.
However, the coverage probability is not effective for our proposed method. This issue arises because we do not implement any calibration method[2] during the general Bayes estimation step. In future work, it will be necessary to include a calibration step.
Table 1: Summary of estimated hazard ratios: The true value of the hazard ratio is 0.607 when the exposure variable increases by 10 units. The sample size is 200, and the number of iterations is 300. The empirical standard error (ESE), percentage bias (%bias), root mean squared error (RMSE), and 95% coverage probability (95%CP) of the estimated hazard ratios for treatment across 300 iterations are summarized for each method in the ‘Method’ column. The ‘Proposed’ method refers to the application of our proposed Bayesian method. The ‘Naive’ and ‘Infeasible’ methods involve the implementation of the ordinary Cox proportional hazards model, without and with the unmeasured covariate, respectively.
Method Summary statistics
ESE %bias RMSE 95%CP
Proposed 0.061 -0.638 0.061 100
Naive 0.022 8.236 0.081 3.67
Infeasible 0.033 -0.005 0.033 93.33
Figure 2: The result of point estimates for each method
5 Discussions and Conclusions
In this manuscript, we introduced a novel nonparametric Bayesian method employing general Bayes techniques to estimate an unbiased hazard ratio in the context of the Cox Proportional Hazards Model (CPHM), particularly when dealing with unmeasured confounders. This method involves a two-phase estimating procedure: 1) identifying clusters based on the likelihood of the exposure variable, and 2) estimating the hazard ratio within each identified cluster.
Implicit in our approach is the assumption that unmeasured confounders exert cluster effects, a necessary consideration for the suitability of our methods. As discussed in Section 2, this assumption is deemed reasonable under the circumstances. Since we cannot directly identify unmeasured variables, our approach of assuming a cluster-based structure is considered a pragmatic solution. However, future work should focus on relaxing this assumption, exploring more flexible approaches to accommodate a broader range of data structures and confounder characteristics.